an archaeological sample has only 25% as much C-14 as in a living sample.Knowing that the half life of C-14 is 5800 years,calculate the age of the sample

Grade:9

1 Answers

Aman Bansal
12 years ago

Dear Sri,

The formula for calculating the age of sample is :

t = {1\over \lambda}\ln {A_o\over A_s}

where Ao/As = 1/4(Since only 25% of living carbon content is left)

lemda = 0.693/5800(half life).

 

Best Of luck

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple  to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert